OpenStudy (anonymous):

Solve -np - 80 > 60 for n. Show your work. - Someone please explain to mee.

OpenStudy (anonymous):

It's just algebraic manipulation like so: \[\eqalign{ &-np-80\gt60 \\ &-np\gt60+80 \\ &-np\gt140 \\ &\frac{-np}{-p}\lt\frac{140}{-p} \\ &n\lt-\frac{140}{p} }\]

OpenStudy (anonymous):

Wow.. You made that look really easy. (x

OpenStudy (anonymous):

Haha I try ;)

OpenStudy (anonymous):

I suppose the trick is that when you have something of the form: \[-a\gt x\] Then: \[a\lt-x\]

OpenStudy (anonymous):

You have to flip the inequality sign around I suppose!

OpenStudy (anonymous):

Noo, what you did is you went about it like this: \[\eqalign{ &-np-80\gt60 \\ &-np\lt60+80 }\] Your mistake is in line two, you do not need to switch the direction of the sign when adding and subtracting, just dividing and multiplying by a negative